Solve y=-x+8 and x+y=7
A) This system has no solutions.
B) This system has infinitely many solutions
C) (1/2, 15/2)
D) (-1/2, 17/2)

Found 2 solutions by stanbon, ewatrrr:
Answer by stanbon(75887)

Solve y=-x+8
and x+y=7
---
Slope forms:
y = -x+8
y = -x+7
---
The slopes are the same and intercepts
are different so the lines are parallel.
Ans: A
--------------
Cheers,
Stan H.
===============
A) This system has no solutions.
B) This system has infinitely many solutions
C) (1/2, 15/2)
D) (-1/2, 17/2)

Answer by ewatrrr(24785)


Hi
Solve
y = -x+8 and x+y=7 OR y = -x +7
Line are parallel, both with slopes = -1
There is no solution (lines have no point in common)